OpenStudy (anonymous):

Which of the following has the least value? Thirty-seven and fourteen hundredths Seventy and thirty-one tenths One hundredth Seventeen and ninety-one thousandths

OpenStudy (lncognlto):

So putting them into an easily comparable form: 37 and 14/100 = 37.14 70 and 31/10 = 73 and 1/10 = 73.1 1/100 = 0.01 17 and 91/1000 = 17.091 Therefore, one hundredth has the least value.